OpenStudy (anonymous):

(50x^(3/4)y/2x^(1/2))^(1/2) + (x^(1/2)y^(1/2)/9x^(3/4)y^(3/2))^(-1/2)

OpenStudy (unklerhaukus):

\[{\left( \frac{50x^{3/4}y}{2x^{1/2}}\right)}^{1/2} + \left(\frac{x^{1/2}y^{1/2}}{9x^{3/4}y^{3/2}}\right)^{-1/2}\]

OpenStudy (anonymous):

yes

OpenStudy (unklerhaukus):

\[={\left( 25{x^{1/4}y}\right)}^{1/2} + \left(\frac{x^{-1/4}y^{-1}}{9}\right)^{-1/2}\]

OpenStudy (unklerhaukus):

\[= 5x^{1/8}y^{1/2} + \left( \frac{1}{9}\frac{1}{x^{1/4}y} \right)^{-1/2}\]

OpenStudy (unklerhaukus):

\[= 5x^{1/8}y^{1/2} + 3{x^{1/8}y^{1/2}}\]

OpenStudy (unklerhaukus):

\[= 8x^{1/8}y^{1/2} \] \[=8\sqrt[8]{x}\sqrt{y}\]

OpenStudy (anonymous):

thank you soo much